OpenStudy (anonymous):

6b + 5 - 6b in simplest form

OpenStudy (anonymous):

I think the 6b would cancel out since theres a positive 6b and a negative.

OpenStudy (blank ):

6b+5-6b According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, b is a factor of both 6b and -6b. (6-6)b+5 To add integers with different signs, subtract their absolute values and give the result the same sign as the integer with the greater absolute value. In this example, subtract the absolute values of 6 and -6 and give the result the same sign as the integer with the greater absolute value. 0+5 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. 5 It looks like 5 is the simplest it could get. :)